OpenStudy (anonymous):

solve the equation x-2/3 = 2(x)/7 the '/' is for a fraction

OpenStudy (anonymous):

x-2/3 = 2x/7 multiply by both 3 and 7 21x-14=6x 14=15x x=14/15